40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16241     Entry time: Thu Jul 8 11:20:38 2021
Author: Anchal, Paco, Gautam 
Type: Summary 
Category: LSC 
Subject: PRFPMI locking attempts 

Last night Gautam walked us through the algorithm used to lock PRFPMI. We tried it several times with the PSL HEPA filter off between 10:00 pm July 7th to 1:00 am July 8th. None of our attempts were successful. In between, we tried to do the locking with old IMC settings as well, but it did not change the result for us. In most attempts, the arms would start to resonate with PRMI with about 200 times the power than without power recycling while the arms are still controlled by ALS beatnote. The handover of lock controls "CARM+DARM locked to ALS beatnote" to "Main laser + IMC locked to the CARM+DARM" would always fail. More specifically, we were seeing that as soon as we hand over the DC control of CARM from ALS beatnote to IR by feeding back to MC2, the lock would inevitably fail before the rest of the high-frequency control can be transferred over.

Nonetheless, Paco and I got a good demo of how to do PRFPMI locking if the need appears. With more practice and attempts, we should be able to achieve the lock at some point in the future. The issues in handover could be due to any of the following:

  • Although it seems like ALS beatnote fed control of arms keep them within the CARM IR linewidth as we see the IR resonating, there still could be some excess noise that needs to be dealt with.
  • Gautam conjectures, that the presence of high power in the arms connects the ITMs and the ETMs with an optical spring changing the transfer function of the pendula. This in turn changes the phase margin and possibly makes the CARM loop in IR PRFPMI unstable.
  • We should also investigate the loop transfer functions near the handover point for the ALS beatnote loop and the IR CARM loop and calculate the crossover frequency and gain/phase margins there.

More insights or suggestions are welcome.


Note; An earthquake came around lunch time and tripped all watchdogs. Most suspensions were recovered without issues, but ITMX appeared to be stuck. We tried the shaking procedure, but after this we couldn't restore the XARM lock. From alignment, we tried optimizing the TRX but we only got up to ~0.5 and ASS wouldn't work as usual. In the end the issue was that we had forgotten to enable the LL coil output devil so after we did this, we managed to recover the XARM.

ELOG V3.1.3-